The final reaction of glycolysis is catalyzed by pyruvate kinase and has a ΔG’° of -31.4kJ/mol. In vivo this enzyme has a ΔG of -16.7kJ/mol. In gluconeogenic tissues like the liver, the cytosolic concentrations of ATP, ADP, Pi and PEP are 2.25mM, 0.25mM, 1.65mM and 8.8μM respectively.

Assuming everything else remains constant, what fold change in pyruvate concentration would be required to drive pyruvate kinase in the reverse direction? (That is, solve the equation when ΔG is 0).
Answer #1

ADP+ Pi = ATP (\DeltaG is positive means it is the reverse process)

\DeltaG0 = \DeltaG0p - \DeltaG0r (summation of products & reactants)

=2.25 - ( 0.25 + 1.65) = 2.25 - 1.90 = 0.35kJ/mol